# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ 1 Introduction
for any $x\in\mathbb{R}^{d}$ and some $\lambda\in\mathbb{R}$, with $\pi$ an unknown function on $\mathbb{R}^{d}$, $\varphi$ an observed forcing on $\mathbb{R}^{d}$ and $\mathrm{k}:\ \mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ a positive integral kernel—where $\pi$ and $k$ belong to suitably regular classes. We focus on the case in which $\pi$ is a positive function with known finite integral, i.e. $\int_{\mathbb{R}^{d}}\pi(x)\mathrm{d}x<\infty$, so that we can renormalize the left hand side of (1) and consider the normalized version of $\pi$ and define the new forcing as $\varphi(x)/\int_{\mathbb{R}^{d}}\pi(x)\mathrm{d}x$ (which we denote by $\pi,\varphi$ too with a slight abuse of notation). These integral equations find applications in numerous contexts including econometrics [20, 12], modelling light transport for ray-tracing and related applications [52, 53] and reinforcement learning [17]. In the homogeneous case (i.e. when $\varphi\equiv 0$), solving (1) corresponds to finding the eigenfunction $\pi$ of $\mathrm{k}$ for a given eigenvalue $\lambda$; the eigen–decomposition of positive kernels $\mathrm{k}$ is normally referred to as Karhunen–Loève decomposition (see [18] for a recent review) and has applications to Gaussian process regression [55], spatial statistics [15, 49] and Markov chain Monte Carlo (MCMC) in infinite dimensional spaces [1]. When $\mathrm{k}$ is the density of a Markov kernel (i.e. $\int\mathrm{k}(x,y)\mathrm{d}y=1$) and $\lambda=1$, solving (1) is equivalent to finding an invariant measure of $\mathrm{k}$. This is useful when dealing with approximate MCMC kernels [2, 44] or tractable approximations to transition densities [7].
@@
where $\textup{KL}\left(\mu\middle|\nu\right)=\int_{\mathbb{R}^{d}}\log((\mathrm{d}\mu/\mathrm{d}\nu)(y))\mathrm{d}\mu(y)$ denotes the Kullback–Leibler divergence between $\mu$ and $\nu$, for a given regularization parameter $\alpha>0$ and some reference measure $\pi_{0}$. Henceforth, we implicitly work under the assumption that the $\pi$ of (1) is the Lebesgue density of a probability distribution over a Euclidean space. Standard approaches normally rely upon discretization of $\pi$ and $\mathrm{k}$ through basis functions or using numerical integration (see, e.g., [5] for a survey of classical methods). These methods are particularly effective in low-dimensional settings, but they become challenging when $d$ is not small and in the case in which the solution $\pi$ is defined over an unbounded interval. In this paper we focus on the latter issue, and develop a method to solve Fredholm integral equations of the second kind on unbounded domains.
@@
We extend the approach of [16] to equations of the second kind and derive a particle approximation to a Stochastic Differential Equation (SDE) of McKean–Vlasov type (an MKVSDE) associated with the Wasserstein gradient flow minimizing (2). This approach is particularly well-suited to solving integral equations on unbounded domains since is not based on a fixed discretization over a finite support, but on an adaptive stochastic discretization with diffusive behaviour.
@@
Finally, we demonstrate empirically that our method provides accurate solutions to integral equations of the form (1) on unbounded domains and can be applied more generally than other methods (e.g. those based on the von Neumann decomposition; see our first numerical experiment). We stress that the presence of the regularization term $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ in (2) is crucial for obtaining significant improvements in the case in which the integral operator induced by the kernel $\mathrm{k}$ is not invertible or poorly conditioned—as one might expect, regularization is critical in order to obtain good solutions to this ill-posed inverse problem.
The remainder of the manuscript is organized as follows. In Section 2 we study the minimization problem in (2) and give conditions under which it admits a unique minimizer. We derive the Wasserstein gradient flow and the corresponding McKean–Vlasov SDE in Section 3. Section 4 introduces an interacting particle system approximating the MKVSDE and a numerical scheme which approximates the continuous time dynamics. We also provide explicit error bounds on the approximation. Finally, in Section 5, we test our method on several examples and compare with other methods commonly used in the literature. Julia code to reproduce all examples is available online at https://github.com/FrancescaCrucinio/FE2kind_WGF.
@@ 1.1 Notation
[…]
@@ 2 The Minimization Problem
We now introduce the minimization problem for the functional (2) and give conditions under which it is well-posed. We then highlight connections with maximum entropy methods, a popular technique to solve Fredholm integral equations in low-dimensional settings.
@@ 2.1 Stability of $\mathcal{F}_{\alpha}$
The density of the kernel $\mathrm{k}$ and the forcing $\varphi$ are such that $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$, $\varphi\in\mathrm{C}^{\infty}(\mathbb{R}^{d},[0,+\infty))$ and there exists $\mathtt{M}\geq 0$ such that for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ we have $\mathrm{k}(x,y)+\left\|\nabla\mathrm{k}(x,y)\right\|+\left\|\nabla^{2}\mathrm{k}(x,y)\right\|\leq\mathtt{M}$ and $\varphi(x)+\left\|\nabla\varphi(x)\right\|+\left\|\nabla^{2}\varphi(x)\right\|\leq\mathtt{M}$.
Assumption 1 restricts the class of kernels $\mathrm{k}$ to smooth kernels with bounded derivatives up to the second order. This might seem restrictive, however, smooth kernels are particularly challenging while discountinuous or degenerate kernels normally allow for specific solution methodologies (see, e.g., [38, Chapter 11]). In this section, we will require only that $\mathrm{k}\in\mathrm{C}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$. The control of higher order derivatives is necessary to establish convergence result for the interacting particle system (both in continuous and discrete time) and will be exploited later.
To address stability issues we also consider the following regularized functional: for any $\alpha,\eta\geq 0$ and $\pi\in\mathcal{P}(\mathbb{R}^{d})$
$\mathcal{F}_{\alpha}^{\eta}(\pi)$ $=\mathcal{F}^{\eta}(\pi)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right),$ (4)
@@
The proof of Proposition 1 shows that the condition $\alpha>0$ is sufficient to guarantee uniqueness of the minimizer, in fact we rely on the coercivity of $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ to show that $\mathcal{F}_{\alpha}^{\eta}$ is coercive. Thus, for the remainder of this work we only consider the case $\alpha,\eta>0$.
@@ 2.2 Maximum Entropy Methods
~$\tilde{\mathcal{F}}_{\alpha}(\pi)=\textup{KL}\left(\pi\middle|\varphi+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)-\alpha\mathrm{H}(\pi),$ (6)
@@
where $\mathcal{P}_{\mathrm{H}}(\mathbb{R}^{d})$ is the set of probability distributions with finite entropy. However, the functional ~$\tilde{\mathcal{F}}_{\alpha}$ is not lower bounded and the corresponding minimization problem is not well-defined and therefore is not considered here.
Maximum entropy solutions to Fredholm integral equations were originally proposed in [43] under moment constraints obtained by selecting different families of basis functions [34, 31]. In the one dimensional case, the maximum entropy solution can be written analytically (see, e.g., [31, Proposition 3.1]), and these approaches are particularly effective. However, in higher dimensions maximum entropy methods generally require discretization of the support of the solution $\pi$.
@@ 3.1 Wasserstein Gradient Flow
To derive the Wasserstein gradient flow PDE for $\mathcal{F}_{\alpha}^{\eta}$ we first derive the Wasserstein subdifferential of the functional [4, Definition 10.1.1], to guarantee the existence of this operator we further assume that $\pi_{0}$ satisfies the following assumption.
@@
    $\pi_{0}$ admits a density w.r.t. the Lebesgue measure, $\mathrm{d}\pi_{0}(x)=\pi_{0}(x)\mathrm{d}x$, with $\pi_{0}(x)\propto\exp\left[-U(x)\right]$, where, $U:\ \mathbb{R}^{d}\to\mathbb{R}$ belongs to $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R})$.
@@
In Appendix B we show that under Assumption 1 and 2(a) the subdifferential of $\mathcal{F}_{\alpha}^{\eta}$ is given by
$\partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi)$ =$=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right.$
@@
The Wasserstein gradient flow [4, Definition 11.1.1] associated with $\mathcal{F}_{\alpha}^{\eta}$ is a family of probability measures $(\pi_{t})_{t\geq 0}$ satisfying (in a weak sense) $\partial_{t}\pi_{t}=\nabla\cdot(\pi_{t}\ \partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi))$, which in our case corresponds to:
$-\nabla\cdot\left(\pi_{t}\left\{\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi_{t}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi_{t}\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi_{t}\left(z\right)-\alpha\nabla U(x)\right\}\right)$ (8)

### v2

@@ 1 Introduction
for any $x\in\mathbb{R}^{d}$ and some $\lambda\in\mathbb{R}$, with $\pi$ an unknown function on $\mathbb{R}^{d}$, $\varphi$ an observed forcing on $\mathbb{R}^{d}$ and $\mathrm{k}:\ \mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ a positive integral kernel—where $\pi$ and $\mathrm{k}$ belong to suitably regular classes. We focus on the case in which $\pi$ is a positive function with known finite integral, i.e. $\int_{\mathbb{R}^{d}}\pi(x)\mathrm{d}x<\infty$, so that we can renormalize the left hand side of (1) and consider the normalized version of $\pi$ and define the new forcing as $\varphi(x)/\int_{\mathbb{R}^{d}}\pi(x)\mathrm{d}x$ (which we denote by $\pi,\varphi$ too with a slight abuse of notation). Such equations find applications throughout statistics and data science in contexts as diverse as econometrics [22, 12], modelling light transport [54, 55] and reinforcement learning [18]. In the homogeneous case (i.e. when $\varphi\equiv 0$), solving (1) corresponds to finding the eigenfunction $\pi$ of $\mathrm{k}$ for a given eigenvalue $\lambda$; the eigen–decomposition of positive kernels $\mathrm{k}$ is normally referred to as Karhunen–Loève decomposition (see [20] for a recent review) and has applications to Gaussian process regression [59], spatial statistics [15, 51] and Markov chain Monte Carlo (MCMC) in infinite dimensional spaces [1]. When $\mathrm{k}$ is the density of a Markov kernel (i.e. $\int\mathrm{k}(x,y)\mathrm{d}y=1$) and $\lambda=1$, solving (1) is equivalent to finding an invariant measure of $\mathrm{k}$. This is useful when dealing with approximate MCMC kernels [2, 46] or tractable approximations to transition densities [7].
@@
where $\textup{KL}\left(\mu\middle|\nu\right)=\int_{\mathbb{R}^{d}}\log((\mathrm{d}\mu/\mathrm{d}\nu)(y))\mathrm{d}\mu(y)$ denotes the Kullback–Leibler divergence between $\mu$ and $\nu$, for a given regularization parameter $\alpha>0$ and reference measure $\pi_{0}$. Henceforth, we implicitly work under the assumption that the $\pi$ of (1) is the Lebesgue density of a probability distribution over a Euclidean space. Standard approaches normally rely upon discretization of $\pi$ and $\mathrm{k}$ through basis functions or using numerical integration (see, e.g., [5] for a survey of classical methods). These methods are particularly effective in low-dimensional settings, but they become challenging when $d$ is not small and in the case in which the solution $\pi$ is defined over an unbounded set. In this paper we focus on the latter issue, and develop a method to solve Fredholm integral equations of the second kind on unbounded domains.
@@
We extend the approach of [16] to equations of the second kind, obtaining a particle approximation of a nonlinear stochastic differential equation (SDE) of McKean–Vlasov type (an MKVSDE) associated with the Wasserstein gradient flow minimizing (2). We also study the convergence of the minimizers of (2) when the regularization parameter $\alpha\to 0$. This approach is particularly well-suited to integral equations on unbounded domains since is not based on a fixed discretization, but on an adaptive stochastic discretization with diffusive behaviour.
@@
Finally, we demonstrate empirically that our method provides good solutions to integral equations of the form (1) on unbounded domains and can be applied more generally than other methods (e.g. those based on the von Neumann decomposition; see our first numerical experiment). We stress that the presence of the regularization term $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ in (2) is crucial for obtaining significant improvements in the case in which the integral operator induced by the kernel $\mathrm{k}$ is not invertible or poorly conditioned—as one might expect, regularization is critical in order to obtain good solutions to this ill-posed inverse problem.
The remainder of the manuscript is organized as follows. In Section 2 we study the minimization problem in (2) and give conditions under which it admits a unique minimizer. We derive the Wasserstein gradient flow and the corresponding McKean–Vlasov SDE in Section 3. Section 4 introduces an interacting particle system approximating the MKVSDE and a numerical scheme which approximates the continuous time dynamics. We also provide explicit error bounds on the approximation. Finally, in Section 5, we test our method on several examples and compare with other methods commonly used in the literature.
@@ 1.1 Notation
We endow $\mathbb{R}^{d}$ with the Borel $\sigma$-field $\mathcal{B}(\mathbb{R}^{d})$ with respect to the Euclidean norm $\left\|\cdot\right\|$; when dealing with a matrix $A$ we consider the spectral norm $\left\|A\right\|:=\sup_{\left\|u\right\|=1}\left\|Au\right\|$ induced by the Euclidean norm of vectors. We denote $\mathrm{C}(\mathbb{R}^{d},\mathbb{R}^{p})$ the set of continuous functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ and $\mathrm{C}^{n}(\mathbb{R}^{d},\mathbb{R}^{p})$ the set of $n$-times differentiable functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ for any $n\in\mathbb{N}^{\star}$. For all differentiable $f$, we denote by $\nabla f$ its gradient. For all $f\in\mathrm{C}^{1}(\mathbb{R}^{d_{1}}\times\dots\times\mathbb{R}^{d_{m}},\mathbb{R}^{p})$ with $m\in\mathbb{N}^{\star}$, we denote by $\nabla_{i}f$ its gradient w.r.t. component $i\in\{1,\ldots,m\}$. Furthermore, if $f$ is twice differentiable we denote by $\nabla^{2}f$ its Hessian and by $\Delta f$ its Laplacian. We say that a function $f:\ \mathsf{X}\to\mathbb{R}$ (where $\mathsf{X}$ is a metric space) is coercive if for every $t\in\mathbb{R}$, $f^{-1}(\left(-\infty,t\right])$ is relatively compact. We denote by $\mathcal{P}(\mathbb{R}^{d})$ the set of probability measures over $\mathcal{B}(\mathbb{R}^{d})$, and endow this space with the topology of weak convergence. For any $p\in\mathbb{N}$, we denote by $\mathcal{P}_{p}(\mathbb{R}^{d})=\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}\left\|x\right\|^{p}\mathrm{d}\pi(x)<+\infty\}$ the set of probability measures over $\mathcal{B}(\mathbb{R}^{d})$ with finite $p$-th moment. We denote the subset of $\mathcal{P}(\mathbb{R}^{d})$ of measures which are absolutely continuous w.r.t. Lebesgue by $\mathcal{P}^{ac}(\mathbb{R}^{d})$ and let $\mathcal{P}^{ac}_{p}(\mathbb{R}^{d}):=\mathcal{P}_{p}(\mathbb{R}^{d})\cap\mathcal{P}^{ac}(\mathbb{R}^{d})$. For any $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ we define the $p$-Wasserstein distance $\mathbf{W}_{p}(\mu,\nu)$ between $\mu$ and $\nu$ by
@@ 2 The Minimization Problem
We first introduce the minimization problem for (2) and give conditions under which it is well-posed, before studying the convergence to the unregularized problem and highlighting connections with maximum entropy methods, a popular technique to solve low-dimensional Fredholm integral equations.
@@ 2.1 Stability of $\mathcal{F}_{\alpha}$
The density of the kernel $\mathrm{k}$ and the forcing $\varphi$ are such that $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$, $\varphi\in\mathrm{C}^{\infty}(\mathbb{R}^{d},[0,+\infty))$ and there exists $\mathtt{M}\geq 0$ such that for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ we have $\mathrm{k}(x,y)+\left\|\nabla\mathrm{k}(x,y)\right\|+\left\|\nabla^{2}\mathrm{k}(x,y)\right\|+\left\|(\partial^{3}_{ij\ell}\mathrm{k}(x,y))_{i,j,\ell}\right\|\leq\mathtt{M}$, where $\partial^{3}_{ij\ell}$ denotes the third derivative of $\mathrm{k}$ w.r.t. component $i,j,\ell\in\{1,2\}$, and $\varphi(x)+\left\|\nabla\varphi(x)\right\|+\left\|\nabla^{2}\varphi(x)\right\|\leq\mathtt{M}$.
Assumption 1 restricts the class of kernels $\mathrm{k}$ to smooth kernels with bounded derivatives up to the third order. This might seem restrictive, however, smooth kernels are particularly challenging while discountinuous or degenerate kernels normally allow for specific solution methodologies (see, e.g., [40, Ch. 11]). In this section, we will require only that $\mathrm{k}\in\mathrm{C}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$. The control of higher order derivatives is used to establish convergence result for the interacting particle system (in both continuous and discrete time) .
To address stability issues we also consider the following regularized functional:
$\forall\alpha,\eta\geq 0,\pi\in\mathcal{P}(\mathbb{R}^{d}):\qquad\mathcal{F}_{\alpha}^{\eta}(\pi)=\mathcal{F}^{\eta}(\pi)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right),$ (4)
@@
The proof of Proposition 1 shows that the condition $\alpha>0$ is sufficient to guarantee uniqueness of the minimizer, in fact we rely on the coercivity of $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ to show that $\mathcal{F}_{\alpha}^{\eta}$ is coercive. For the remainder of this work we only consider the case $\alpha,\eta>0$.
2.2 Convergence of Minimizers
We now consider the limiting behaviour of the regularized functional $\mathcal{F}_{\alpha}^{\eta}$ when the parameters $\alpha,\eta$ converge to 0. In particular, we show that if both parameters tend to 0 the minimizer of $\mathcal{F}_{\alpha}^{\eta}$ converges to one of the minimizers of the unregularized functional.
To control the behaviour of the family of minima $\inf_{\pi\in\mathcal{P}^{ac}(\mathbb{R}^{d})}\{\mathcal{F}_{\alpha}^{\eta}(\pi):\ \alpha,\eta>0\}$ when $\alpha$ is close to 0, we make use of the following assumption which controls the tail behaviour of $\pi_{0}$ and the tail behaviour of the density $\mathrm{k}$.
Assumption 2.
The following hold:
  1. (a)
[…]
    $C_{1}-\tau\left\|x\right\|^{2}\leq U(x)\leq C_{1}+\tau\left\|x\right\|^{2}.$
  2. (b)
    there exists $C_{2}\geq 0$ such that for any $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$
    $\mathrm{k}(x,y)\geq C_{2}^{-1}\exp[-C_{2}(1+\left\|x\right\|^{2}+\left\|y\right\|^{2})].$
Under this assumption we have the following results which describe the behaviour of the minimizer as $\eta\to 0$ and as $(\alpha,\eta)\to(0,0)$ jointly.
Proposition 2.
Assume 1 and 2.
  1. 1.
    If there exist $\pi_{\alpha}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}}$ such that $\lim_{n\to+\infty}\eta_{n}=0$ and $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{\alpha,\eta_{n}}^{\star},\pi_{\alpha}^{\star})=0$, then
    $\pi_{\alpha}^{\star}\in\arg\min_{\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})}\mathcal{F}_{\alpha}(\pi).$
  2. 2.
    If there exist $\pi^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}},(\alpha_{n})_{n\in\mathbb{N}^{\star}}\in(0,+\infty)^{\mathbb{N}}$ such that $\lim_{n\to+\infty}\eta_{n}=0,\lim_{n\to+\infty}\alpha_{n}=0$ and $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{\alpha,\eta_{n}}^{\star},\pi^{\star})=0$, then
    $\pi^{\star}\in\arg\min_{\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})}\mathcal{F}(\pi).$
Proof.
See Appendix B. ∎
@@ 2.3 Maximum Entropy Methods
$\tilde{\mathcal{F}}_{\alpha}(\pi)=\textup{KL}\left(\pi\middle|\varphi+\lambda\int_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\pi(y)\mathrm{d}y\right)-\alpha\mathrm{H}(\pi),$ (6)
@@
where $\mathcal{P}_{\mathrm{H}}(\mathbb{R}^{d})$ is the set of probability distributions with finite entropy. However, $\tilde{\mathcal{F}}_{\alpha}$ is not lower bounded and the corresponding minimization problem is not well-defined and therefore is not considered here.
Maximum entropy solutions to Fredholm integral equations were originally proposed in [45] under moment constraints obtained by selecting different families of basis functions [36, 33]. In the one dimensional case, such a solution can be written analytically (see, e.g., [33, Prop. 3.1]), and these approaches are particularly effective. However, in higher dimensions maximum entropy methods generally require discretization of the support of the solution $\pi$.
@@ 3.1 Wasserstein Gradient Flow
To derive the Wasserstein gradient flow PDE for $\mathcal{F}_{\alpha}^{\eta}$ we first derive the Wasserstein subdifferential of the functional [4, Def. 10.1.1], to guarantee the existence of this operator we further assume that $\pi_{0}$ satisfies the following assumption.
@@
[…]
@@
In Appendix C we show that under Assumption 1 and 3(a) the subdifferential of $\mathcal{F}_{\alpha}^{\eta}$ is:
$\partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi)$ $=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right$.
@@
The Wasserstein gradient flow [4, Def. 11.1.1] associated with $\mathcal{F}_{\alpha}^{\eta}$ is a family of probability measures $(\pi_{t})_{t\geq 0}$ satisfying (in a weak sense) $\partial_{t}\pi_{t}=\nabla\cdot(\pi_{t}\ \partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi))$, which in our case corresponds to:
$\partial_{t}\pi_{t}=$ (8)
$-\nabla\cdot\left(\pi_{t}\left\{\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi_{t}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi_{t}\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi_{t}\left(z\right)-\alpha\nabla U(x)\right\}\right)$